OpenStudy (yamato_delgado):

How to solve this questions? "A fence is to be built to enclose a rectangular area of 800 square meter. The fence along three sides is to be made of material that costs 6 dollars per meter, and the material for the fourth side costs 18 dollars per meter. Find the dimensions of the rectangular that will allow the most economical fence to be built. What is the most economical dimension if the cost of all four side of the fence is the same? "

OpenStudy (sooobored):

First, we start by defining some variables let x be the width of our rectangle let y be the length of our rectangle 

OpenStudy (sooobored):

Let "C" be the total cost of the fencing material forgot to include, variables x and y are in meters Then, we look at the given information and attempt to place them into the form of an equation "A fence is to be built to enclose a rectangular area of 800 square meter" means 800= x*y since the area of a rectangle is given by A= w*l "The fence along three sides is to be made of material that costs 6 dollars per meter, and the material for the fourth side costs 18 dollars " this means that 2x + 1y side are 6 dollars per meter and the remaining side y is 18 dollars per meter. we can represent this as 6(2x+y) is equal to the total cost of the three sides and 18y is the total cost of the 4th side and putting them together to get the total cost of the fencing material "C" C= 6(2x+y)+18y

OpenStudy (sooobored):

So now we look at what the question is asking us. "What is the most economical dimension if the cost of all four side of the fence is the same? " in other words, what are the dimensions of the rectangle in order to get the lowest cost C

OpenStudy (sooobored):

but wait, C is a "function" of two variables x and y we cant do a single derivative in order to determine a critical point--- well you could, but i dont think it would help much since there is a relationship between x and y

OpenStudy (sooobored):

Thankfully, we remember that the area of the rectangle is also a function of x an y this means we can solve for either x or y and then substitute that into the main cost "C" equation so x=800/y or y=800/x hooray

OpenStudy (sooobored):

now, since this is an optimization problem where we're looking for the most economical cost, lets assume that the graph of the length of one side vs total cost will look something like this